OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. Write your solution in set brackets with no spaces, such as {-1,2}. 2x^2-28x+98=0

OpenStudy (anonymous):

2x^2 - 28x + 98 = 0 x^2 - 14x + 49 = 0 (x-7)^2 = 0 x = 7

OpenStudy (anonymous):

brackets ?

OpenStudy (anonymous):

{7}